If 0.5 mole of BaCl2 is mixed with0.2 mole of Na3PO4 the maximum number of mole Ba3(PO4)2 that can be formed is?
1
Expert's answer
2013-12-02T07:23:07-0500
Reaction:
3BaCl2 + 2Na3PO4 = Ba3(PO4)2 + 6NaCl

The mole ratio between BaCl2 and Na3PO4 is
n(BaCl2):n(Na3PO4)=3:2

The ratio of these compounds in the given mixture is 0.5/0.2 = 5:2 > 3:2.
Hence the BaCl2 is in excess and the Na3PO4 is the limiting reactant. The 1 mole of Ba3(PO4)2 occurs for each two moles of Na3PO4, hence the amount of Ba3(PO4)2

n[Ba3(PO4)2] = 1/2 * 0.2 = 0.1 mol.